help please
rome File Edit View History Bookmarks People Window Help ■Chapter24egnment G consider The Three StociX + x 0BA2720-107 2018/2019xID created by Camtasia StoX s彫Catnal pacebook y Twitter D SMS 慝Amazon wSJ Poptropica @ C&IB Relationship-muriedin Help Save &Exit Submit hapter 2 Assignment Check my work 3 Consider the three stocks in the at time t Stock C splits two-for-one in the last period following table. Pe represents price at time t and oe represents shares outstanding 166 points 100 82 200 2 200 4 400 a. Calculate the rate of return on a price-weighted index of the three stocks for the first period(t-O to 1 (Do not round intermediate calculations. Round your answer to 2 decimal places) Print b. What will be the divisor for the price weighted i your answer to 2 decimal places index in year 2? (Do not round intermediate calculations. Round < Pree 3016 Next > esc 20 15
Answer #1

Solution(s):

a)  Calculate the rate of return on a price-weighted index of the three stocks for the first period (t=0 to t=1)


At t = 0,

Price value of the index = (81+41+82)/3

Price value of the index = 68


At t = 1,

Price value of the index = (86+36+92)/3

Price value of the index = 71.33


The rate of return = (71.33/68) - 1

The rate of return = 4.90%


(b)What must happen to the divisor for the price-weighted index in year 2?


In the absence of a split, stock C would sell for 110, and the value of the index would be:
(86+36+92)/3 = 71.33

After the split, stock C sells at 55. Therefore, we need to set the divisor (d) such that:
71.33 = (86+36+46)/d
d = 168/71.33

d= 2.36

Divisor for the price-weighted index in year 2 = 2.36